OpenStudy (anonymous):

Find the following by using definition: v=t - 1/t I don't know what definition is =3=

OpenStudy (anonymous):

?

OpenStudy (amistre64):

i think it means that the equation is defined as v = t - 1/t

OpenStudy (anonymous):

Well I'm not sure what the definition of definition is but I'm guessing you have to differentiate it using a certain method which I don't really know of.

OpenStudy (amistre64):

are you trying to find the derivative ?

OpenStudy (amistre64):

D[v = t - 1/t] D[v] = D[t - 1/t] D[v] = D[t] - D[1/t]

OpenStudy (amistre64):

im not sure what route your spose to take as is; do you have a screen shot or picture that we can parse?

OpenStudy (anonymous):

Exactly the question was 'Find the following indicated derivatives by using definition' this is just one of the questions given.

OpenStudy (anonymous):

haha I just realized that I didn't say 'find the derivative' in the original question. My bad :D

OpenStudy (amistre64):

well, first principles is a pain with limits and whatnots; derivative rules are much easier

OpenStudy (amistre64):

the rules are defined as: D[k] = 0, k is any constant D[x^n] = n x^(n-1) D[f+g] = D[f]+D[g] but im not sure if thats what is meant by "definition"

OpenStudy (anonymous):

Any idea on what it'd look like by using the first principle method?

OpenStudy (beginnersmind):

I'm pretty sure they want you to find the derivative by using the definition of the derivative. So you can't used the power rule. The definition would be: \[v'(t)=\lim_{h \rightarrow 0}\frac{v(t+h)-v(t)}{h}\]

OpenStudy (anonymous):

I've tried that and had difficulty in solving it :(

OpenStudy (beginnersmind):

I agree, it looks messy. Can you post the question word by word?

OpenStudy (amistre64):

how is this thing graded, is it a written work, or do you simply need to submit a final result?

OpenStudy (beginnersmind):

@amistre64 LOL

OpenStudy (amistre64):

:)

OpenStudy (anonymous):

I'm just studying for an upcoming exam ^.^ and exactly the question looks like this: \[v=t-1/t\] I don't know how to present it in any other form. I don't get the LOL...

OpenStudy (amistre64):